sinB = cos(B-45), where B and (B-45) are acute angles.Find the value of the angle B?

Grade:10

2 Answers

Arun
5 years ago
Dear student
 
By allied angles conversion
90 – B = B – 45
 
2B = 135
 
hence B = 67.5 degree
 
hope it helps
 
Regards
Arun
Rishit Dagli
5 years ago
We know, 
If sin a = cos b
Then 90-a=b
Therefore,
90-b=b-45
2b=135
 
Hope it helps 
Rishit Dagli
b=135/2
b=67.5 degrees